OpenStudy (anonymous):

how do you find the derivative of (y+4)/(x-2)=5y at (3,1)???

OpenStudy (amistre64):

either solve for "y" all by its little lonesome; or use implicit deriv..different.....that one word...

OpenStudy (amistre64):

lets try implicit... 5y = (y+4)/(x-2) ; take the d/dx of both sides.. (d/dx)(5y) = (d/dx)[(y+4)/(x-2)]

OpenStudy (amistre64):

5y' = apply quotient rule to this side 5y' = [(x-2)(1y') - (y+4)(x')] / [(x-2)^2] ; dy/dx = y' ; x' = 1

OpenStudy (amistre64):

get y' by itself if you dare :)

OpenStudy (anonymous):

no easier way? how can you get y by it self

OpenStudy (amistre64):

lol.... I get: y+4 y' = ------ x-7

OpenStudy (anonymous):

how??

OpenStudy (amistre64):

plug in (3,1) y' = -5/6 ?? where did I lose ya? :)

OpenStudy (amistre64):

y' = -5/4 ; forgot how to add lol

OpenStudy (anonymous):

how did you get to y+4/x-7

OpenStudy (amistre64):

by using paper and pencil really... but lets see if I can type it out...

OpenStudy (amistre64):

(x-2)(1y') - (y+4)(x') 5y' = ---------------- ; move (x-2)^2 up and bring y' down.. (x-2)^2 (x-2)(y') - (y+4) 5(x-2)^2 = ---------------- ; split the fraction.... y' (x-2)(y') (y+4) 5(x-2)^2 = ------- - ------- ; cancel top to bottom your y' y' y' (y+4) 5(x-2)^2 = (x-2) - ------- ; now subtract (x-2) from both sides y' - (y+4) [5(x-2)^2]-(x-2) = ------- ; I forgot something on the paper.... y'

OpenStudy (amistre64):

regardless tho, lets work this out: -4-y y' = ---------------- [5(x-2)^2] - (x-2) try that one :)

OpenStudy (amistre64):

(x=3,y=1) y' = (-4-1) / [5(3-2)^2 - (3-2)] y' = -5 / 4 , is that right?

OpenStudy (amistre64):

HAH!!! I got the same answer as I did with the last one lol

OpenStudy (anonymous):

yessss thank you so much